OpenStudy (anonymous):

HELP I GOT THE WRONG ANSWER Medal+Fan If f(x) = 3x + 5 and g(x) = 2x - 9, find f(x) + g(x). A. 5x + 4 B. x - 4 C. x + 14 D. 5x - 4 I choose C as the answer but i got it wrong

OpenStudy (demandandsupply):

Make sure you keep the (2x-9) in brackets

OpenStudy (demandandsupply):

so you have 5-9, not 5+9

OpenStudy (mathmale):

You might want to re-arrange your two functions into columns, as shown: f(x) = 3x + 5 + g(x) = 2x - 9 ---------------- Simply add up each of the 3 columns shown. Write your answer as f(x)+g(x) = ?x + ?

OpenStudy (anonymous):

f(x) = 3x + 5=x + g(x) = 2x - 9=4.5 @mathmale

OpenStudy (mathmale):

Simply add up each of the 3 columns shown. f(x) = 3x + 5 + g(x) = 2x - 9 ---------------- f(x)+g(x) = 5x - 4 Note that we're NOT trying to solve for x, so x=4.5 is not what is called for. We're just combining the two linear functions to obtain f(x)+g(x) = 5x - 4.

OpenStudy (mathmale):

@Ambbiiee : Are you OK with the solution of this problem, or would you like more discussion of it?